what is the x=y+8 5x+3y+12 i dont get systems of linear equations using substitution ive tryed x=y+8 8+x -y but i mean i still dont get it
Answer by mananth(16946)

x=y+8
OR y=x-8.................(1)
5x+3y=12 ................(2)
5x+3y=12 .............1
y=-8+x ---------2
Substitute y in (1)
5x+3(-8+x)=12
5x-24+3x=12
5x+3x=12+24
8x=36
/ 8
x= 4.5
Plug the value of x in (1)
5x+3y=12
5*4.5+3y=12
22.5+3y =12
3y=-10.5
/3
y=-3.5
m.ananth@hotmail.ca